# Anomalous Spontaneous Emission Enhancement by Non-Hermitian Momentum-Space Bound States in the Continuum

### v1

@@ Abstract
Conventional Purcell theory emphasizes high quality factors ($Q$) for spontaneous emission (SE) enhancement in cavities, but overlooks collective Bloch mode effects in periodic nanostructures like photonic crystal slabs. We introduce a unified temporal coupled-mode framework to compute Purcell and photoluminescence factors through momentum-space integration, revealing anomalous SE enhancement by non-Hermitian momentum-space bound states in the continuum (BICs). In silicon gratings with comparable effective mode volumes, this yields substantial SE enhancement in low-$Q$ regimes—defying the traditional high-$Q$ paradigm and inversely correlated with system QQ—while emission rates are stably twice the photoluminescence, eliminating critical coupling requirements. Unique spectral profiles, contradicting Lorentzian/Fano assumptions, arise from collective mode interactions. Full-wave simulations confirm these challenges to conventional wisdom, with non-Hermitian BICs outperforming high-$Q$ designs across broad numerical apertures. This establishes a novel paradigm leveraging non-Hermiticity and topological protection for robust, bright emitters, redefining nanophotonic applications in lasers and light-emitting diodes.
@@
The enhancement of spontaneous emission (SE) in optical cavities has long been a cornerstone of photonics. Traditionally, the Purcell factor establishes a direct connection between the cavity quality factor ($Q$) and the emission rate[1], motivating decades of efforts toward high-$Q$ resonators[2, 3], such as plasmonic cavities[4, 3] and photonic crystal defect modes[5, 6, 7].
@@
This distinction is central to two complementary quantities: the local density of states (LDOS, i.e., the Purcell factor)[14, 15], which accounts for total emission power from the full momentum space, and the photoluminescence (PL) factor[16, 17, 18, 19, 20, 21, 22], which reflects the energy collected in the far field and thus only samples a restricted subset of momentum states. Two theoretical frameworks have been developed in parallel—LDOS based on Fermi’s golden rule[12, 23, 24], and PL based on local field enhancement[25, 26, 27], linked through Lorentz reciprocity[28]. Yet, existing theories often consider only specific wavevectors[29, 30] and neglect the collective role of Bloch modes, leaving an incomplete picture of light–matter interactions in periodic structures. A unified framework specialized for periodic systems is still lacking.
@@
Here, we demonstrate that this collective effect yields phenomena that challenge conventional intuition. We first develop a generalized temporal coupled-mode theory (TCMT)[31, 32], grounded in input-output theory[33], for multi-mode, non-Hermitian systems. This framework enables simultaneous evaluation of the Purcell factor ($F_{p}$) and the photoluminescence enhancement factor ($F_{\text{PL}}$). Applying this framework to a minimal yet common two-level system, we quantify the collective effect by integrating emission power across momentum space. Surprisingly, we find an anomalous SE enhancement near BICs, which is negatively correlated with the system’s overall $Q$ factor, contradicting the conventional Purcell formula. We validate our prediction through grating designs that achieve substantial SE enhancement with regular mode volume, unattainable with conventional design approaches. Furthermore, we find that the total emission power is consistently twice the PL power, i.e., $F_{p}\approx 2F_{\text{PL}}$, indicating the absence of critical coupling between radiative and non-radiative channels, which relaxes fabrication tolerances and enhances design flexibility. The frequency-domain lineshape is also unique, exhibiting neither Lorentzian nor Fano characteristics. Finally, our framework encompasses both BICs and QGMs, providing precise guidance for designing and optimizing high-performance nanophotonic emitters in versatile periodic systems.
@@
Figure 1: Schematic of SE enhancement in local and nonlocal metasurfaces. (a) Local metasurface with absorbing boundaries (e.g., PML), where dipoles couple to discrete modes via QNMs. (b) Nonlocal metasurface with periodic boundaries, supporting orthogonal Bloch modes labeled by in-plane wavevector $k$. (c) Continuous band structure in momentum space, where a dipole excites all modes at a given frequency. (d) PL collected from the radiative power $P_{\rm rad}$ within a limited NA in momentum space.
@@
As shown in Figure 1, SE enhancement differs fundamentally between local and nonlocal metasurfaces. In local systems (Figure 1(a)), a dipole couples to a few discrete modes, typically described by quasinormal modes (QNMs)[34, 12]. In contrast, periodic metasurfaces (Figure 1(b)) support orthogonal Bloch modes across momentum space (Figure 1(c))[14], so a dipole at fixed frequency excites infinitely many $k$-states, enabled by periodic Bloch boundary conditions, unlike the open boundaries of local systems[12]. Consequently, the LDOS, or Purcell factor ($F_{p}$), is evaluated by integrating the total emission power $P_{\rm tot}$ over the entire momentum space (Figure 1(c))[35]. The PL factor ($F_{\text{PL}}$), however, is determined by integrating the far-field radiation power $P_{\rm rad}$ within a restricted momentum space limited by the numerical aperture (NA) of the objective lens (Figure 1(d))[30].
@@
Physically, Eq. (1) isolates far-field PL, Eq. (2) captures non-radiative loss, and Eq. (3) recovers the conventional LDOS[35, 36, 37]. Their ratio, $\eta(\omega)=\frac{P_{\rm rad}(\omega)}{P_{\rm tot}(\omega)}$, gives the radiation efficiency, equivalent to the external quantum efficiency (EQE)[38]. Full derivations are provided in the Supplementary Materials Sec. 1[39]. This compact formulation enables clear decomposition of energy flows. When absorption- and radiation-related losses are comparable, this leads to phenomena such as critical coupling[26]. Particularly when interacting with ultrahigh-$Q$ resonances (e.g., BICs), this separation is essential, even in low-loss dielectrics[40]. We note that similar energy-partition treatments have recently been discussed for gain–loss systems[24]; in contrast, our formulation targets passive structures and PL observables, providing a directly applicable tool for nanophotonic emitters.
@@
We implement a ternary Brillouin-zone folding (BZF)[41] in a silicon grating (Figure 2(a)) to construct a well-defined pair of orthogonal TE-polarized BIC and QGM modes at the $\Gamma$ point, with their field profiles shown in Figure 2(c). The ternary BZF is realized by periodically modulating the widths of every three grating ridges by the factor $S$, while maintaining a constant filling factor ($\text{DC}=0.6$) and thickness ($t=a/3$, where $a$ is the lattice constant), as illustrated in Figure 2(b). For comparison, binary folding produces different effects, as discussed in the Supplementary Materials Sec. 5.[39] The silicon grating has a refractive index of 3.5, and to capture realistic non-radiative losses, we introduce an imaginary part of $10^{-3}$ to the refractive index. By varying $S$, we effectively control the system’s non-Hermiticity: the symmetry-protected BIC exhibits vanishing radiation loss, while the QGM acquires radiative decay $\gamma\propto S^{2}$ (Figure 2(e)) and the detuning between modes scales as $\delta\propto S$ (see Supplementary Materials Sec. 5 for details[39])[42].
@@
where the integrals sum over all Bloch modes in momentum space. Three approximations are implicit here. First, the emitter acts as a localized point source with a finite spatial extent, resulting in a radiation pattern that limits effective coupling to high-$k$ Bloch modes and thus imposes finite integration bounds. Second, PL collection optics have a limited NA, imposing a cutoff $|k|<{\rm NA}\cdot k_{0}$, where $k_{0}$ is the free-space wavevector. More rigorous consideration is to set $F_{\rm PL}(\omega)\propto{\int_{-{\rm NA}\cdot k_{0}}^{{\rm NA}\cdot k_{0}}P_{\rm rad}(\omega,k)\,dk}/{(2{\rm NA}\cdot k_{0})}$, as elaborated in Supplementary Materials Sec. 2.[39] Third, far-field measurements are performed in angular coordinates, requiring $d\theta$ rather than $dk$.Nevertheless, this approach adequately captures the collective effects and aligns well with full-wave simulations, so we proceed without these refinements here. Further details on these approximations are provided in Supplementary Materials Sec. 2.[39]
@@
Building on these theoretical predictions, we validate them using full-wave simulations. As indicated in Figure 3(a, e), we model a one-dimensional silicon grating with refractive indices $n_{\text{high}}=3.5+10^{-3}i$ and $n_{\text{low}}=\Re[n_{\text{low}}]+10^{-3}i$, analyzing the behavior of TM modes (Figure 3(b, f)). The grating maintains a constant filling factor ($\text{DC}=0.6$) and thickness ($t=a/3$). The effective Hamiltonian in Eq. (4) remains sufficiently applicable to this system. By tuning the refractive index contrast ($\Re[n_{\text{high}}]-\Re[n_{\text{low}}]$), we control the system’s non-Hermiticity. In the strongly non-Hermitian regime ($\Re[n_{\text{low}}]=1$), the BIC remains robust despite a sharp decrease in its $Q$ factor under symmetry breaking, exhibiting flat dispersion around the BIC[44], as reflected in the featured far-field PL spectra (Figure 3(h)). In contrast, a grating with weak non-Hermiticity ($\Re[n_{\text{low}}]=3$) shows a higher $Q$ factor and intrinsic dispersion (Figure 3(c, d)). Notably, despite the large difference in non-Hermiticity between the two systems, the spatial field distributions of the BICs remain nearly identical (Figure 3(b, f)). This implies similar effective mode volumes $V_{\text{eff}}$, which isolates the $Q$ factor’s role in SE enhancement.

### v2

@@ Abstract
Conventional Purcell theory emphasizes high quality factors ($Q$) for spontaneous emission (SE) enhancement in cavities, but overlooks collective Bloch mode effects in periodic nanostructures like photonic crystal slabs. We introduce a unified temporal coupled-mode framework to evaluate Purcell and photoluminescence factors through momentum-space integration, revealing anomalous SE enhancement by non-Hermitian momentum-space bound states in the continuum (BICs). In periodic dielectric structures with comparable effective mode volumes, this yields substantial SE enhancement in low-$Q$ regimes—defying the traditional high-$Q$ paradigm and inversely correlated with system QQ—while emission rates are stably twice the photoluminescence, eliminating critical coupling requirements. Unique spectral profiles, contradicting Lorentzian/Fano assumptions, arise from collective mode interactions. Full-wave simulations confirm these challenges to conventional wisdom, with non-Hermitian BICs outperforming high-$Q$ designs across broad numerical apertures. This establishes a novel paradigm leveraging non-Hermiticity and topological protection for robust, bright emitters, redefining nanophotonic applications in lasers and light-emitting diodes.
@@
The enhancement of spontaneous emission (SE) in optical cavities has long been a cornerstone of photonics. Traditionally, the Purcell factor establishes a direct connection between the cavity quality factor ($Q$) and the emission rate[1], motivating decades of efforts toward high-$Q$ resonators[2, 3, 4], such as plasmonic cavities[5, 3] and photonic crystal defect modes[6, 7, 8].
@@
This distinction is central to two complementary quantities: the local density of states (LDOS)[15, 16], which accounts for the total emission power, giving rise to the Purcell factor via comparison with the free-space emission rate, and the photoluminescence (PL) factor[17, 18, 19, 20, 21, 22, 23], which reflects the radiation energy collected in the far field. Two theoretical frameworks have been developed in parallel—LDOS based on Fermi’s golden rule[13, 24, 25], and PL based on local field enhancement[26, 27, 28], linked through Lorentz reciprocity[29]. Yet, existing theories often consider only specific wavevectors[30, 31] and neglect the collective role of Bloch modes, leaving an incomplete picture of light–matter interactions in periodic structures. A unified framework specialized for periodic systems is still lacking.
@@
Here, we demonstrate that this collective effect yields phenomena that challenge conventional intuition. We first develop a generalized temporal coupled-mode theory (TCMT)[32, 33], grounded in input-output theory[34], for multi-mode, non-Hermitian systems. This framework enables simultaneous evaluation of the Purcell factor ($F_{p}$) and the photoluminescence enhancement factor ($F_{\text{PL}}$). Applying this framework to a minimal yet common two-level system, we quantify the collective effect by integrating emission power across momentum space. Surprisingly, we find an anomalous SE enhancement near BICs, which is negatively correlated with the Q factors at every wavevector, contradicting the conventional Purcell formula. We validate our prediction through grating designs that achieve substantial SE enhancement with regular mode volume, unattainable with conventional design approaches. Furthermore, we find that the total emission power is consistently twice the PL power, i.e., $F_{p}\approx 2F_{\text{PL}}$, indicating the absence of critical coupling between radiative and non-radiative channels, which relaxes fabrication tolerances and enhances design flexibility. The frequency-domain lineshape is also unique, exhibiting neither Lorentzian nor Fano characteristics. Finally, our framework encompasses both BICs and QGMs, providing precise guidance for designing and optimizing high-performance nanophotonic emitters in versatile periodic systems.
@@
Figure 1: Schematic of SE enhancement in local and nonlocal metasurfaces. (a) Local metasurface with absorbing boundaries (e.g., PML), where dipoles couple to discrete modes. (b) Nonlocal metasurface with periodic boundaries, supporting orthogonal Bloch modes labeled by in-plane wavevector $\mathbf{k}$. (c) Continuous band structure in momentum space, where a dipole excites all modes at a given frequency. (d) PL collected from the radiative power $P_{\rm rad}$ within a limited NA in momentum space.
@@
As shown in Figure 1, SE enhancement differs fundamentally between local and nonlocal systems. In local systems with open boundaries (Figure 1(a))[13], a dipole couples to modes[35, 13] characterized by discrete order $n$. In contrast, the periodic Bloch boundary conditions introduce the in-plane wavevector $\mathbf{k}$ as another continuous label, thereby enabling periodic metasurfaces (Figure 1(b)) to support orthogonal Bloch modes across momentum space (Figure 1(c))[15] and allowing a dipole at fixed frequency to excite infinitely many k-states.[36] Consequently, the LDOS, or Purcell factor ($F_{p}$), is evaluated by integrating the total emission power $P_{\rm tot}$ over the entire momentum space (Figure 1(c)). The PL factor ($F_{\text{PL}}$), however, is determined by integrating the far-field radiation power $P_{\rm rad}$ within a restricted momentum space limited by the numerical aperture (NA) of the objective lens (Figure 1(d))[31].
@@
Physically, Eq. (1) isolates far-field PL, Eq. (2) captures non-radiative loss, and Eq. (3) recovers the conventional LDOS[37, 38, 39]. Their ratio, $\eta(\omega)=\frac{P_{\rm rad}(\omega)}{P_{\rm tot}(\omega)}$, gives the radiation efficiency, equivalent to the external quantum efficiency (EQE)[40]. Full derivations, along with an alternative perspective using QNMs, are provided in Supplementary Materials Sec. 1[41]. This compact formulation enables clear decomposition of energy flows. When absorption- and radiation-related losses are comparable, this leads to phenomena such as critical coupling[27]. Particularly when interacting with ultrahigh-$Q$ resonances (e.g., BICs), this separation is essential, even in low-loss dielectrics[42]. We note that similar energy-partition treatments have recently been discussed for gain–loss systems[25]; in contrast, our formulation targets passive structures and PL observables, providing a directly applicable tool for nanophotonic emitters.
@@
We implement a ternary Brillouin-zone folding (BZF)[43] in a silicon grating (Figure 2(a, b)) to construct a well-defined pair of orthogonal TE-polarized BIC and QGM modes at the $\Gamma$ point, with their field profiles shown in Figure 2(c). The ternary BZF is realized by periodically modulating the widths of every three grating ridges by the factor $S$, while maintaining a constant filling factor ($\text{DC}=0.6$) and thickness ($t=a/3$, where $a$ is the lattice constant), as illustrated in Figure 2(b). For comparison, binary folding produces different effects, as discussed in the Supplementary Materials Sec. 5.[41] The silicon grating has a refractive index of 3.5, and to capture realistic non-radiative losses, we introduce an imaginary part of $10^{-3}$ to the refractive index. By varying $S$, we effectively control the system’s non-Hermiticity: the symmetry-protected BIC exhibits vanishing radiation loss, while the QGM acquires radiative decay $\gamma\propto S^{2}$ (Figure 2(e)) and the detuning between modes scales as $\delta\propto S$ (see Supplementary Materials Sec. 5 for details[41])[44].
@@
where the integrals sum over all Bloch modes in momentum space. Three approximations are implicit here. First, for a non-ideal point source emitter, the finite size limits the effective coupling to high-$k$ modes, thus imposing finite integration bounds. Second, PL collection optics have a limited NA, imposing a cutoff $|k|<{\rm NA}\cdot k_{0}$, where $k_{0}$ is the free-space wavevector. More rigorous consideration is to set $F_{\rm PL}(\omega)\propto{\int_{-{\rm NA}\cdot k_{0}}^{{\rm NA}\cdot k_{0}}P_{\rm rad}(\omega,k)\,dk}/{(2{\rm NA}\cdot k_{0})}$, as elaborated in Supplementary Materials Sec. 2.[41] Third, far-field measurements are performed in angular coordinates, requiring $d\theta$ rather than $dk$.Nevertheless, this approach adequately captures the collective effects and aligns well with full-wave simulations, so we proceed without these refinements here. Further details on these approximations are provided in Supplementary Materials Sec. 2.[41]
@@
Building on these theoretical predictions, we validate them using full-wave simulations. As indicated in Figure 3(a, e), we model a one-dimensional silicon grating with refractive indices $n_{\text{high}}=3.5+10^{-3}i$ and $n_{\text{low}}=\Re[n_{\text{low}}]+10^{-3}i$, analyzing the behavior of TM modes (Figure 3(b, f)). The grating maintains a constant filling factor ($\text{DC}=0.6$) and thickness ($t=a/3$). The effective Hamiltonian in Eq. (4) remains sufficiently applicable to this system. By tuning the refractive index contrast ($\Re[n_{\text{high}}]-\Re[n_{\text{low}}]$), we control the system’s non-Hermiticity. In the strongly non-Hermitian regime ($\Re[n_{\text{low}}]=1$), the BIC remains robust despite a sharp decrease in its $Q$ factor under symmetry breaking, exhibiting flat dispersion around the BIC[46](Figure 3(g)), as reflected in the featured far-field PL spectra (Figure 3(h)). In contrast, a grating with weak non-Hermiticity ($\Re[n_{\text{low}}]=3$) shows a higher $Q$ factor and intrinsic dispersion (Figure 3(c, d)). Notably, despite the large difference in non-Hermiticity between the two systems, the spatial field distributions of the BICs remain nearly identical (Figure 3(b, f)). This implies similar effective mode volumes $V_{\text{eff}}$, which isolates the $Q$ factor’s role in SE enhancement.